phoebeR:

Which of the following is a polynomial with roots negative square root of 3 , square root of 3, and 2?

XioGonz:

Answer choices?

phoebeR:

x3 + 2x2 − 3x − 6 x3 − 2x2 − 3x + 6 x3 + 3x2 − 5x − 15 x3 − 3x2 − 5x + 15

XioGonz:

Thank you.

phoebeR:

how do I solve it?

XioGonz:

You have to use the FOIL method, are you familiar with it?

phoebeR:

yes

XioGonz:

Alright, what is your first step

phoebeR:

I don't know

XioGonz:

1 attachment
phoebeR:

how do i do that with the square roots?

XioGonz:

First, multiply first terms of each binomial: q*q=q2

XioGonz:

Here, I'll just give you the answer shhh x^3-2x^2-3x+6.

phoebeR:

there aren't any binomials in this, how did you get that answer?

phoebeR:

hello, can you please explain how to solve this?

darkknight:

Which of the following is a polynomial with roots negative square root of 3 , square root of 3, and 2? \[-\sqrt(3), \sqrt(3), 2\] are the roots So \[(x+\sqrt(3))( x-\sqrt(3))(x-2)\]

darkknight:

And then expand from there

darkknight:

\[(x+\sqrt(3))(x-\sqrt(3)) = x^2 + xsqrt(3)-xsqrt(3) -3 = x^2-3\] and then now \[(x^2-3)(x-2)\]

darkknight:

You good now?

phoebeR:

I think so thank you!

darkknight:

np, please click best response, up to you though. Have a good day